OpenStudy (anonymous):

rico's first three test scores in biology were 65,90, and 73.what was his mean score.

OpenStudy (anonymous):

65+90+73 = 228 divided by three = 76

OpenStudy (anonymous):

The mean is given by the sum over the components divided by the total number of components...

OpenStudy (anonymous):

to find the mean, add all the values and divide by the number of them... \[\frac{ 65+90+73 }{ 3 } =??\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw...

OpenStudy (anonymous):

how are u guys doing